OpenStudy (anonymous):

a professional football team has a stadium which seats 60000. it is found that x tickets can be sold at a price of p=10-(x/10000) dollars per ticket. find the value of x and p at which the total money received will be a maximum.

OpenStudy (amistre64):

thats a linear equation with a positive slope....

OpenStudy (amistre64):

errr.. a negative slope lol;

OpenStudy (amistre64):

p = -(1/100000) x + 10

OpenStudy (amistre64):

when x = 0 you have a maximum of 10; the higher x goes, the less you make

OpenStudy (amistre64):

or am i missing something

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I don't think so. There's logic in your answer ^_^

OpenStudy (amistre64):

this might be the derivative and is asking fo ryouto integrate from 0 to 60000

OpenStudy (anonymous):

hmmm, no wait, wouldn't that be a minimum?

OpenStudy (anonymous):

if 0 tickets are sold, then you gained p=10000

OpenStudy (amistre64):

f'(x) = 0 at 10000/10 then

OpenStudy (anonymous):

oh right lol, it's over (10,000) ^_^" my bad lol

OpenStudy (amistre64):

f'(x)= 1000; integrate from 0 to 60000?

OpenStudy (anonymous):

no need to integrate, from the function , you can figure out the maximum, just like you did :)

OpenStudy (anonymous):

x increases= less money, like you've said. so the maximum is at x = 0?

OpenStudy (amistre64):

well i am thinking that this is the derivative of a higher function; and its max value is at 1000 for x; so the integral would tell us the price right?

OpenStudy (anonymous):

but wouldn't wouldn't the price be found after plugging in x =0 in the function? :)

OpenStudy (amistre64):

if this was the true function then yes, but the wording has me believe that the optimazation is given by this equation; and not the true function...

OpenStudy (anonymous):

mmm, so we have to optimize in this case? ^_^to find the maximum

OpenStudy (amistre64):

i think so :) so we would integrate the given to find the true value; with x = 1000; becasue that is where p'(x) = 0

OpenStudy (anonymous):

Alright, now I got what you mean lol proceed ^_^

OpenStudy (amistre64):

the 60000 gives us an initial condition; possiply at (0,60000)

OpenStudy (anonymous):

so we'll integrate from 0->6000?

OpenStudy (amistre64):

dunno if we integrate of simply anchor it in at (0,60000) p=10-(x/10000)

OpenStudy (amistre64):

P(x)=10x -x^2/2(10000) + C

OpenStudy (amistre64):

10(1000) - (1000)(1000)/2(10000) 10000 - 500/10

OpenStudy (amistre64):

10000 - 50 = 9950 +C

OpenStudy (amistre64):

P(1000) = 60000 = 9950 + C C = 60000 - 9950 = 50050 P(x) = 10(x) - x^2/20000 + 50050 ??

OpenStudy (amistre64):

maybe integrate from 0 to 60000 :) ??? dunno

OpenStudy (anonymous):

I think we have made it look more complicated than it seems, lol?

OpenStudy (anonymous):

LOL!

OpenStudy (anonymous):

do you know what I think?

OpenStudy (amistre64):

broken question? lol

OpenStudy (anonymous):

since the maximum price of tickets is achieved when x=0, then for 60000 seats, the maximum price will be = 600,000

OpenStudy (anonymous):

.-. right?

OpenStudy (amistre64):

maximum price = 600,000? or max revenue?

OpenStudy (amistre64):

see the wording is off...

OpenStudy (anonymous):

hmm, we have :\[p(x) = 10-\frac{x}{10000}\] so if we take x = 0: \[p(0) = 10\] so we can tell that the original price of the ticket is = 10. Now for 60000 seats, we need 60000 people, so each person is going to buy a ticket. In the end, maximum price achieved = 600,000 I'm not sure, but I'm just computing what I have understood from the question ^_^" ahaha...

OpenStudy (anonymous):

so the maximum values for x and p : x = 0, and p = 10

OpenStudy (amistre64):

i agree that that is a possible interpretation of the problem :) but maximum price is not 600,000; that would be revenue. Max price if anything would be 10 lol

OpenStudy (anonymous):

oh, right >_< well, atleast you got what I mean =P

OpenStudy (dumbcow):

x=50,000 p=5 maximum money = p*x = 250,000

OpenStudy (amistre64):

if x = tickets sold, then how can 0 ticket sold bring in 600,000 :)

OpenStudy (anonymous):

where did you get the 50,000 and 5? .-.

OpenStudy (dumbcow):

money received is revenue which equals price*quantity quantity is num seats sold = x so the function to maximized is p*x = x(10-x/10,000)

OpenStudy (anonymous):

annnnnnwaaaar~ lol ^_^

OpenStudy (anonymous):

but the number of seats = 60,000 and not 50k lol :)

OpenStudy (anonymous):

I don't know where you guys have reached. But, I think the idea here is to maximize the total money received. Total money= price of one ticket * # of tickets= xp(x).

OpenStudy (dumbcow):

60,000 available not sold

OpenStudy (anonymous):

so how did you get 50k?

OpenStudy (amistre64):

i understand it now.... i think lol

OpenStudy (dumbcow):

differentiate above function set =0 solve for x

OpenStudy (anonymous):

p = 0 and solve for x?

OpenStudy (amistre64):

whoever wrote this problem could have been clearer ;)

OpenStudy (anonymous):

xp(x)=f(x)=10x-x^2/10,000 f'(x)=10-2x/10,000=0 implies x=50,000

OpenStudy (anonymous):

oh well, clear or not, my brain is shutdown and doesn't want to study. I had calculus II at 8 and was playing in the whole time, and got out not understanding what on Earth I took today.

OpenStudy (anonymous):

and I felt good abt it. I guess it's beacuse of the bananas I ate in the morning.

OpenStudy (anonymous):

oh, I got you anwar, thank you ^_^

OpenStudy (anonymous):

LOL!!

OpenStudy (anonymous):

You're welcome. Am I right there?

OpenStudy (anonymous):

it's TRUE! Bananas = get you excited, and I was off ~ completely OFF

OpenStudy (anonymous):

hoenstly, I have no idea. My brain is unable to compute or accept any info atm.

OpenStudy (anonymous):

but it looks right though ^_^

OpenStudy (amistre64):

i was watching the cheerleaders and got distracted lol

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

HAHA.

OpenStudy (dumbcow):

waterlily do you get it? or do you need more help

OpenStudy (anonymous):

my professor tried to grab my attention by trying, "TRYING" to introduce Taylor series today, but no hope, he gave up and ended the class after 1 hour =D!

OpenStudy (anonymous):

I was like, I don't like taylor series, I want to learn Fourier series!

OpenStudy (anonymous):

he was like, but it's not in the syllabus and etc etc etc

OpenStudy (anonymous):

i get it thanks

OpenStudy (anonymous):

loooooooooooong morning, it was HILARIOUS

OpenStudy (anonymous):

oh, sorry waterlily ^_^" aha...

OpenStudy (anonymous):

You have no more classes?

OpenStudy (anonymous):

let's see, my ESPII professor, cancelled today's class at 10, so no class til 3:30 , hallalujah~!

OpenStudy (anonymous):

and guess what!

OpenStudy (anonymous):

what?!

OpenStudy (anonymous):

I've got a test at that time! Heck knows what I'll during that hour. If I'm luney and off now, what will happen later on? ^_^

OpenStudy (anonymous):

I think I and mr. "dumbcow" had the same answer. He was first though.

OpenStudy (anonymous):

people usually revise, take advantage of these 4 free hours, but me? nooooooo, don't feel like it ._. I guess I need a day off.

OpenStudy (anonymous):

lol.....LOL!

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

why did he call himself dumbcow?! :D

OpenStudy (anonymous):

I find that a really funny and "creative" username :)

OpenStudy (anonymous):

I have no idea.

OpenStudy (anonymous):

cows are not dumb! they just don't get english that's all

OpenStudy (anonymous):

Do you stay at university or go home during this 4 hours?

OpenStudy (anonymous):

if cows are dumb for not understanding what we say, then we are dumb becase we don't understand what they say, right?

OpenStudy (anonymous):

LOL. they don't get english heh? what about arabic?

OpenStudy (anonymous):

I stay here, I don't want to get my mum tired.

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

all they say is mooo

OpenStudy (anonymous):

I don't agree with your "cow" theorem btw :P

OpenStudy (anonymous):

with different wavelengths ~

OpenStudy (anonymous):

But I like it :)

OpenStudy (anonymous):

Hey, this is logic and I'm studying it this semester and I agree with my theorem

OpenStudy (anonymous):

omg..I sound drunk LOL with bananas~

OpenStudy (anonymous):

Nope, it's not.